The school BETA club is looking over the money it collected from its last dance. Faculty tickets cost half as much as student tickets. Student tickets purchased at the door sold for twice as much as the tickets purchased ahead of time. Twice as many students bought tickets at the door as they did in advance, and the total number of students who bought in advanced was three times as many as tickets purchased by faculty. If the total amount of money collected was $775, for 250 tickets, how much did a student pay in advance?

Answer by mananth(16949)

student tickets --------- $2x each
Faculty ticket ---------------$x each
Student tickets at the door = 4x
No of student tickets in advance =y
No of student tickets at the door =2y

Number of tickets by faculty = y/3
Total tickets = 10y/3 =250
10y=750
y=75

y*2x+2y*4x+(y/3)*x=775
150x+600x+25x=775
775x=775
x=$1
Student ticket bought in advance =$2